OpenStudy (marcelie):

consider the parametric equations. x=4cost y=5sint -pi/2

OpenStudy (marcelie):

eliminate the parametric to find Cartesian equation of the curve i got x^2/16 + y^2 /25 = 1 but how would i graph it with the arrows ?

satellite73 (satellite73):

start at \(-\frac{\pi}{2}\) or rather that point

OpenStudy (marcelie):

oh when you graph it like the direction of the curve

satellite73 (satellite73):

which i think is \((0,-5)\)

OpenStudy (marcelie):



satellite73 (satellite73):

then go counter clockwise

satellite73 (satellite73):

not sure you go all the way around though

OpenStudy (marcelie):

hmm how you know its counterclockwise ?

satellite73 (satellite73):

cause that is the positive direction, and you are going from \(-\frac{\pi}{2}\) to \(\frac{\pi}{2}\) HOWEVER

satellite73 (satellite73):

i think you only get half the ellipse

OpenStudy (marcelie):

ohh

satellite73 (satellite73):

actually now that i look at it that was a lousy answer i gave you know it goes that way because as \(t\) increases from \(-\frac{\pi}{2}\) to \(\frac{\pi}{2}\) , \(y=4\sin(t)\) increases from \(-4\) to \(4\)

satellite73 (satellite73):

ok from -5 to 5 really but you get the idea

satellite73 (satellite73):

and \(x=4\cos(t)\) goes from 0 to 1 and back to 0

satellite73 (satellite73):

but you don't go all the way around the ellipse, since you only have half the period for sine and cosine

OpenStudy (marcelie):

ohh k